Two angles are supplementary. One is 10degrees more than four times the other. Find the measures of the angles.​ (Supplementary angles are angles whose sum is 180degrees​.)
Answer by ikleyn(52793)

.

Let x be the measure of the other angle.


Then the measure of the "one" angle is  4x+10.


You have this equation


    x + (4x+10) = 180  degrees.


    5x = 180-10

    5x = 170

     x = 170/5 = 34.


ANSWER.  The angles are  34 degrees and  180-34 = 146 degrees.